# TimeSeriesBench: An Industrial-Grade Benchmark for Time Series Anomaly Detection Models

### v1

@@ 3.4. Algorithms
We adopt 17 semi-supervised/unsupervised methods that have raised significant discussion and wide-ranging impact in the field of univariate time series anomaly detection. The algorithms and implementation details are listed in Appendix B.5. From the perspective of method attributes, these methods can be divided into statistical methods and deep learning methods.
@@
Statistical methods. These methods, including Sub-LOF (Breunig et al., 2000), SAND (Boniol et al., 2021), and MatrixProfile (Zhu et al., 2018), directly compute the minimum similarity between the current and the historical time window. This manner, however, would make the model highly sensitive to the data characteristics in specific scenarios, resulting in a lack of generalizability.
@@
Prediction-based. These methods aim to predict the ”normal” value of the current metric according to the adjacent observations. If the ground truth is far from the predicted value, an anomaly alarm will be triggered.
Reconstruction-based. These methods aim to denoise the anomaly points by the encoding-decoding phase. If the reconstructed value is far from the ground truth, it can be assumed that an anomaly occurs.
VAE-based. These methods believe that the distribution mapping and sampling process can enhance the denoising effect of the model and learn a more robust representation of normal patterns.
General Time Series Model. The designed universal model structure in this class can be used for all time series tasks including time series forecasting, classification, and anomaly detection. They tend to introduce a more general inductive bias that aims to better represent the general features of time series.
Figure 6. Overall performance ranking using different learning schema under reduced-length $F1_{best}$. Methods are ranked in descending order according to the average of all scores. The row names denote the names of methods, while the column names denote the names of the datasets. The best score of each column is underlined. Due to the inherent assumptions behind the statistical methods being incompatible with the all-in-one and zero-shot schemas, these methods are not included in the ranking, and the related score cells are set to grey without values.
4. Experimental Results
@@
4.1. Overall Performance
@@
The overall ranking list is presented in Fig. 6, including the performance of each method using naive, all-in-one, and zero-shot schema, respectively. Since statistical methods, especially those based on streaming, heavily rely on the premise of data being identically distributed, these methods are not well-suited for the all-in-one and zero-shot learning schemas. Therefore, we only evaluate them under the naive schema. First and foremost, (RI 1) The performance of various models under the naive and all-in-one learning schema can exhibit substantial disparities. It is thought-provoking that these differences do not show a clear bias, suggesting that these variations are the result of a combination of multifaceted factors, which is discussed in detail in Sec. 4.2. Additionally, (RI 2) given the consideration of errors arising from random dataset partitioning, the models demonstrated a relatively consistent performance across the all-in-one and zero-shot modes, as there may exist interdependence among time series in the same dataset.
(RI 3) Generally, the statistical methods exhibit great performance on the dataset with relatively stable shapelets and low noise level (UCR), but work poor on other datasets with more noise. For deep learning based models, contrary to our expectations, the overall performances of the models that contain more complicated structures are not satisfactory. According to our examination across existing datasets, a substantial portion of anomalies in real-world datasets exhibit noticeable deviations from their very close neighboring points (sudden spikes or drops), especially for AIOPS, WSD, and Yahoo. These anomalies can be categorized as either contextual outliers or global outliers, both of them belonging to point-wise anomaly type. Due to its ability to capture very local temporal correlations and its ease of convergence, AR surpasses a number of modern networks in cases where point-wise anomalies dominate. We present a case (Fig. 7) from the AIOPS dataset to vividly demonstrate this phenomenon. Even though these point-wise anomalies seem to be quite evident, a significant portion of the methods are unable to handle such events effectively, most likely because they are highly sensitive to noise and may struggle to generalize well. More cases are listed in Appendix D. As numerous factors can lead to a decrease in performance for DNNs, (RI 4) we summarize the major factors that result in the underperformance of some deep learning methods in point-wise anomaly detection tasks:
High noise level. As shown in the former case, the noise can introduce additional variability and random fluctuations in the data, making it harder for the model to distinguish between genuine anomalies and noise-induced fluctuations.
Lack of training data. Models with complicated structures usually require a larger volume of training data to avoid overfitting due to their higher degrees of freedom and flexibility. Similarly, models that employ intricate loss functions often necessitate a greater amount of training data, including diverse examples, to prevent the emergence of falling into local minima and even trivial solutions.
Trends with large periods. Some data may exhibit long-term trends or patterns that span large periods. These trend components can cause the data distribution to deviate from the comfort zone of the model, causing the model to calculate anomaly scores based on a biased data distribution it has never encountered before.
Unreasonable inductive bias. Deep learning models rely on their architectural design and optimization algorithms to detect anomalies in a semi-supervised manner. If the chosen model architecture or optimization approach is not well-suited for point-wise anomaly detection, the models may have an unreasonable inductive bias that hampers their ability to accurately detect anomalies.
@@
Figure 7. Illustrations that some methods fail to effectively detect point-wise anomalies in certain scenarios. The black line represents the original curve, while the red line represents the anomaly scores provided by the method. Anomalous segments are highlighted in pink.
To demonstrate the impact of (RI 4) in a more impressive manner, we conduct an intriguing experiment under the all-in-one learning schema, where the training data is abundant and diverse. In this experiment, we apply first-order differencing to all the training data before feeding it into the model for training. Subsequently, we evaluate the model’s performance on the same first-order differenced dataset. Despite the potential increase in noise level caused by first-order differencing on the original data, it significantly magnify point-wise anomalies while alleviating the effect brought by the continuous and consistent trend factor, as is vividly shown in Fig. 8. Benefiting from this, some methods can achieve a performance improvement of more than 30% on this dataset after the differencing process. It is important to note that we do not recommend relying solely on differenced data for anomaly detection. Differencing can make the temporal pattern information more ambiguous and noisy, significantly reducing the model’s ability to detect pattern-wise anomalies and thus lowering the detection ceiling. More importantly, it is imperative for newly proposed methods to place greater emphasis on these factors to avoid poor results on these distinct point-wise anomalies.
Figure 8. Illustrations that first-order differencing can amplify the degree of outlierness of point-wise anomalies while introducing noise with higher variance.
4.2. Performance in Fine-Grained Scenarios
@@
Merely referring to aggregated data at the granularity of datasets is insufficient to gain a profound understanding of the model’s performance. We carry out a fine-grained comparison of model performance from the perspectives of both learning schema and anomaly type. The UCR dataset is widely acknowledged for its comprehensive annotation of various anomaly types, making it a well-labeled dataset. Additionally, each time series in the UCR dataset is guaranteed to contain only one anomaly, which facilitates the categorization of different anomaly types. To further refine our analysis, we partition the UCR dataset into two subsets based on its supplemental materials (Wu, 2022b): one subset exclusively consisting of point-wise anomalies and another subset focusing solely on pattern-wise anomalies. We form a 2x2 matrix to reflect the variations in model performance under different conditions by incorporating the dimensions of ”naive” and ”all-in-one” learning schemas, along with different types of anomalies, as shown in Fig. 9.
@@
Figure 9. The 2x2 matrix reflecting the variations in model performance under different conditions. We denote prediction, reconstruction, VAE-based, and general time series methods in light red, yellow, green, and blue backgrounds.
@@
From the boxplots, (RI 5) no solution can always outperform the others in all situations. It is also apparent that (RI 6) the majority of models exhibit a significantly higher detection performance for point-wise anomalies compared to pattern-wise anomalies regardless of the learning schema employed. On a more detailed level, (RI 7) prediction-based models excel in identifying point-wise anomalies, as the steep peak/valley is unpredictable in most cases. Due to the inclusion of a more diverse data distribution under all-in-one learning schema, the models are more likely to learn robust representations of patterns. (RI 8) As a result, under the all-in-one learning schema, there is a significant improvement in the detection performance of point-wise anomalies. This observation is particularly pronounced in the case of the Yahoo dataset each time series of which is relatively short. With regard to pattern-wise anomalies which are considered to be more challenging, the situation becomes more complex and interesting. (RI 9) Methods that aim to generate the temporal window after projecting them to low-dimensional representations (Donut and FCVAE) surpass all others in detecting pattern-wise anomalies, while the performance greatly declines when under the all-in-one mode. The former indicates that the assumption of low-dimensional representations hardly reconstructing high-level pattern-wise anomaly indeed works. An illustration is presented to confirm this in Fig. 9(a), where Donut can easily handle the ”smooth” high-level anomaly. However, the performance deteriorates or even becomes ineffective under the all-in-one mode (Fig. 9(b)). We hypothesize that the capacity of low-rank representations may struggle to cover the diverse data distributions present in various curves. Also, (RI 10) although the general time series methods (TimesNet, OFA, FITS) perform well in other time series tasks, they struggle to outperform methods specifically designed for anomaly detection tasks. This may indicate that there is a gap between time series anomaly detection tasks and general time representation tasks. For example, anomaly detection tasks may require models to have a stronger denoising effect. These gaps need special attention when designing models for anomaly detection tasks.
@@
Figure 10. Case study on CIMIS44AirTemperature2 which contains an easily understandable pattern-wise anomaly.
4.3. Trade-off Between Performance, Cost, and Efficiency
When deploying algorithms in real-world scenarios, it is often necessary to strike a balance between performance, storage costs, and inference speed. Notably different from other surveys, in this benchmark, we prefer inference time over training time due to the practical needs of anomaly detection. Each factor plays a critical role in determining the feasibility and practicality of the deployed algorithm. The trade-offs between these considerations are important to ensure an efficient and effective deployment that meets the specific requirements and constraints of the application. For example, real-time monitoring or critical systems may prioritize efficiency to detect anomalies promptly and respond quickly, while IoT (Internet of Things) devices with limited on-chip memory have to focus on the performance of models with a small number of parameters. We exclude the statistical methods, as their inference time is not a fixed value (polynomially correlated with the volume of historical data). Also, We exclude TFAD because it does not align with the real-time manner. As shown in Fig. 11, the inference time for a single sample is far less than 50 milliseconds for all methods under our experimental platform. The learning-based AR exhibits the most favorable gain-to-cost ratio among all methods. FCVAE takes the longest time to inference due to its multi-epoch MCMC process. EncDecAD has the second longest inference time because it utilizes LSTM to perform inference for 100 time steps. (RI 11) In addition to providing guidance for practical applications, we aim for this perspective and toolkit to assist in discovering the scaling law in the field of time series anomaly detection under more reasonable inductive bias and larger parameter spaces.
Figure 11. The triad of the model’s performance, cost, and efficiency trade-off. The x-axis represents the total inference time (batch size set to 1) of the detectors on AIOPS_7103fa0f-cac4-314f-addc-866190247439 (around 140,000 samples) under default parameters. The y-axis represents the average performance of the models on all datasets under the all-in-one learning schema. The size of the scatter points denotes each method’s cube root of the number of parameters.
5. Conclusion AND FUTURE DIRECTIONS
In this paper, we propose TimeSeriesBench, a comprehensive and application-oriented benchmark for evaluating the performance of existing and emerging UTS anomaly detection methods. TimeSeriesBench takes into account existing industrial concerns and conducts a comprehensive performance evaluation of some well-known and latest methods under settings that meet industrial requirements. Also, it offers unprecedented perspectives for measuring algorithm performance, meanwhile laying a solid foundation for the development of Foundation Models in the field. Moreover, TimeSeriesBench includes a user-friendly toolkit and leaderboard, offering easy-to-use interfaces that allow researchers and practitioners to focus on advancing their algorithms without getting entangled in repetitive tasks. We intend to incorporate more latest methods/evaluation protocols and employ more high-quality data. Also, we will continue to monitor the latest developments in general foundation models for time series within the realm of anomaly detection, adapting them accordingly within our benchmark. Additionally, whether you propose deep learning or statistical methods, we welcome your participation in our leaderboard to help drive the development of the time series anomaly detection community.
Appendix A Delay Constraint
We cope with the issue c (in Sec. 3.3) by imposing strict constraints on the detection latency. As depicted in the illustration (Fig. 12), assuming the latency limit ($k$) is set to 3, an anomaly is considered effectively detected only if it is identified within three sampling points after its occurrence. We designate this strategy as k-delay adjustment. This measure enables a more precise assessment of whether the model can meet the requirement of the scenario where there is a high demand for real-time responsiveness. It is equally essential to acknowledge that this approach is applicable only to datasets whose anomalies are labeled without positional bias. We conduct experiments on the selected datasets with non-biased labels. The overall results in shown in Table 2.
Figure 12. Illustrations of k-delay adjustment. This strategy can be combined with point-adjustment as a complementary evaluation paradigm.
Appendix B Details about Experimental settings
B.1. Experimental Platform
B.2. Datasets
B.3. Learning Schema
B.4. Evaluation Metrics
Following the event-based evaluation criterion outlined in Sec. 3.3, we establish two concrete metrics, namely $F1_{best}$ and $AUPRC$. $F1_{best}$ denotes the highest F1 score calculated under reduced-length point-adjustment when iterating over all possible thresholds. $AUPRC$ calculates the area under the precision-recall curve generated according to reduced-length point-adjustment, which is widely applied in cases where class imbalance is present in the data. While $F1_{best}$ can measure the best detection performance that the model can achieve on the current test set, $AUPRC$ provides a more nuanced evaluation of the model’s performance across different levels of recall, which can be important in anomaly detection. A model with a higher $AUPRC$ tends to be more robust.
B.5. Details of Algorithms
B.5.1. Descriptions of Algorithms
Here is the publication date and detailed introduction of all baselines. We will continuously incorporate new methods, and if you are willing to integrate your approach into our benchmark, please feel free to initiate a pull request on GitHub or send us an email.
MatrixProfile (2018) (Zhu et al., 2018). MatrixProfile is a data structure with corresponding algorithms that can measure the minimum distance between the current and the historical time series subsequence, and detect if there is an anomaly according to the measurement.
SAND (2021) (Boniol et al., 2021). SAND leverages the K-shape (Paparrizos and Gravano, 2016) to cluster normal patterns and compares the upcoming time series subsequence with the cluster centroid in a streaming manner.
Sub-LOF (2000) (Breunig et al., 2000). Sub-LOF is an unsupervised learning algorithm that identifies outliers by measuring the local deviation of a given data point with respect to its neighbors.
AR (2005) (Rousseeuw and Leroy, 2005). AutoRegression (AR) holds the assumption that there exists a linear dependence between each individual instance and its past observations. We implement the model using a linear layer in Pytorch and apply first-order differencing on all the data to eliminate potential trend components as the AR model requires time series data to be stationary.
LSTMAD (2015) (Malhotra et al., 2015). LSTMAD employs Long Short-Term Memory (LSTM) networks to capture the temporal dependency among the time windows. As the paper does not provide a detailed description of the algorithm and there is no official source code available, we implement LSTMAD in two forms: LSTMAD-$\alpha$ in a seq2seq manner and LSTMAD-$\beta$ in a multi-step prediction manner.
AE (2011) (Ng et al., 2011). AutoEncoder (AE) projects a time window that ends at the current timestamp to the lower-dimensional latent space and subsequently reconstructs the window. The reconstructed value of the last timestamp is compared with the ground truth value to calculate the anomaly score.
EncDecAD (2016) (Malhotra et al., 2016). EncDecAD additionally introduces LSTM to AE to obtain more complicated representations.
Donut (2018) (Xu et al., 2018). Donut utilizes variational autoencoder and modified ELBO loss to denoise the anomalies and further learn the robust representation of normal patterns.
FCVAE (2024) (Wang et al., 2024). Frequency-enhanced Conditional Variational Autoencoder (FCVAE) incorporates both global and local frequency characteristics into the conditioning of the CVAE framework, greatly improving the accuracy of reconstructing the normal data.
SRCNN (2019) (Ren et al., 2019). SRCNN borrows the technique of visual saliency detection, combining Spectral Residual (SR) and Convolutional Neural Network (CNN) together to detect time series anomalies.
AnomalyTransformer (2022) (Xu et al., 2022). AnomalyTransformer utilizes an attention mechanism to compute the association discrepancy.
TFAD (2022) (Zhang et al., 2022b). TFAD exploits the information in both time and frequency domains to amplify the traits while creating various data augmentation methods to overcome the lack of anomaly data.
TranAD (2022) (Tuli et al., 2022). TranAD incorporates the principles of adversarial learning to develop a training framework with two stages while integrating the strengths of self-attention encoders to capture the temporal dependency embedded in the time series.
TimesNet (2023) (Wu et al., 2023). TimesNet is a general time series model that is capable of handling multiple tasks such as time series forecasting, anomaly detection, and classification. It folds the time series to 2-D space according to frequency information and uses a vision backbone to capture the 2-D dependencies within the reshaped time series.
OFA (2023) (Tian Zhou, 2023). One-Fits-All (OFA) finetunes the pre-trained GPT backbone on time series data to address the issue of lacking a large amount of data for training.
FITS (2024) (Anonymous, 2024a). FITS employs only a single layer frequency domain linear transformation to model the general temporal dependencies and achieves commendable results on general time series tasks with a minimal number of training parameters.
B.5.2. Implementation details of Algorithms
Appendix C TimeSeriesBench Toolkit
The overview of the toolkit is shown in Fig. 13. The dataset interface and algorithm interfaces are provided to allow evaluations on new or private datasets and novel methods under all learning schemas. All plots of raw curves and scores are saved for specific investigation. Previously mentioned flaws in evaluation criteria have sparked a research fervor and several latest works are dedicated to providing evaluation protocols according to different assumptions (Kim et al., 2022; Wagner et al., 2023; Huet et al., 2022; Paparrizos et al., 2022a; Doshi et al., 2022; Scharwächter and Müller, 2020). Thus we also expose an interface for swiftly developing neo-criteria for evaluation grounded in realistic assumptions and evaluating the performance of all methods w.r.t various datasets. In some scenarios, there might be a trade-off between model accuracy and model training/testing cost. Therefore, we also provide a runtime statistics interface to conveniently track and analyze the runtime information of the model.
Figure 13. Overview of TimeSeriesBench Toolkit.
We provide four kinds of workflows to meet different research needs for time series anomaly detection: algorithms benchmarking, algorithm development, evaluation criteria development, and performance analysis. The following are typical audiences for TimeSeriesBench toolkit:
  • Algorithm Researches. TimeSeriesBench offers a flexible interface for implementing, training, and testing new algorithms. Researchers can choose all the mentioned schemas for model training and testing. The framework also provides a full pipeline for loading datasets, running experiments, conducting evaluations, and performing analysis such as generating plots and comparing anomaly scores.
  • Evaluation Researchers. TimeSeriesBench offers a flexible interface for implementing evaluation protocols based on anomaly scores and ground truth labels. Researchers can easily evaluate existing methods according to their specific protocols. The framework also allows evaluation based on offline scores of methods generated by merely once training and test phase.
  • Practitioners of Community or Enterprise. TimeSeriesBench provides a unified and clear dataset format, making it easy to introduce private datasets. It also allows for easy performance comparison of baselines on custom datasets, providing overall performance in CSV format based on protocols suitable for specific applications. Moreover, TimeSeriesBench allows practitioners to record runtime statistics such as model parameter size and inference time, enabling trade-offs between performance, cost, and efficiency.
Compared to existing benchmark suites on univariate time series anomaly detection (Jacob et al., 2020; Lai et al., 2021; Paparrizos et al., 2022b; Wenig et al., 2022), our benchmark suite incorporates real-world scenarios and challenges, aiming to replicate the complexities and nuances encountered in practical applications. This enables researchers and practitioners to gain deeper insights into the strengths and weaknesses of various deep learning approaches, and is well-prepared for the emergence and development of Foundation Models (FMs) in the field of time series anomaly detection. For more information about TimeSeriesBench, please refer to our project and the website of the leaderboard.
Table 1. Comparison among TimeSeriesBench and existing UTS anomaly detection benchmark suites. TimeSeriesBench presents more angles for performance evaluation, meanwhile offering user-friendly interfaces for up-to-date model and evaluation protocol developments.
Suites Exathlon(Jacob et al., 2020) TODS(Lai et al., 2021) TSB(Paparrizos et al., 2022b) TimeEval(Wenig et al., 2022) Ours
Data Source Real-World
Synthetic
Learning Schema one-by-one
all-in-one
zero-shot
Supported Eval-Criteria point-based
range-based
event-based
Extensibility Dataset
Method
Eval-Criteria
RT statistics
Appendix D Informative Cases on Point-wise Anomalies
Other cases are presented below to demonstrate how the lack of training data (Fig. 14) or trends (Fig. 15) with large periods induce poor performance for point-wise anomalies. Some methods are unable to yield meaningful results on most time series, which may result from their unreasonable inductive bias.
Figure 14. Case study on Yahoo_A3Benchmark-TS13 shows that lack of training data induces poor performance.
Figure 15. Case study on Yahoo_A3Benchmark-TS14 shows that trends with large periods induce poor performance. The scores are obtained under the all-in-one learning schema, eliminating the impact of insufficient data.
Appendix E Supplementary statistics
E.1. Overall Performance
Detailed experimental results under all learning schemas are listed in Table 3,  4,  5. Best scores are highlighted in bold, and second-best scores are highlighted in bold and underlined. Due to the inherent assumptions behind the statistical methods being incompatible with the all-in-one and zero-shot schemas, the related scores are set to ”-”.
E.2. Impacts of Different Evaluation Criteria
We also provide the evaluation results under point-wise PA (Point Adjustment), reduced-length PA, and event-wise PA respectively (Table 6,  7,  8). Point-wise PA often results in inflated scores, while event-wise PA completely neglects the length of the anomaly segments. Best scores are highlighted in bold, and second-best scores are highlighted in bold and underlined. Due to the inherent assumptions behind the statistical methods being incompatible with the all-in-one and zero-shot schemas, the related score are set to ”-”.
E.3. Impacts of Different Learning Schemas
Table  9 shows the performance differences on all datasets using different learning schemas. Best scores are highlighted in bold, and second-best scores are highlighted in bold and underlined. Performance improvements exceeding 5% are marked with $\uparrow$, while performance declines exceeding 5% are marked with $\downarrow$.
E.4. Performance, Cost, Efficiency Trade-off
The model’s performance, cost, and efficiency trade-off under naive and zero-shot learning schema are shown in Fig. 16 and Fig. 17.
Figure 16. The triad of the model’s performance, cost, and efficiency trade-off. The y-axis represents the average performance of the models on all datasets under the naive learning schema. The size of the scatter points denotes each method’s cube root of the number of parameters.
Figure 17. The triad of the model’s performance, cost, and efficiency trade-off. The y-axis represents the average performance of the models on all datasets under the zero-shot learning schema. The size of the scatter points denotes each method’s cube root of the number of parameters.
Table 2. Performance comparison using all-in-one learning schema under K-delay reduced-length $F1_{best}$ and $AUPRC$

### v3

@@ III-D Algorithms
We adopt 17 semi-supervised/unsupervised methods that have raised significant discussion and wide-ranging impact in the field of univariate time series anomaly detection. From the perspective of method attributes, these methods can be divided into statistical methods and deep learning methods.
@@
Statistical methods. These methods, including Sub-LOF [41], SAND [42], and MatrixProfile [43], directly compute the minimum similarity between the current and the historical time window. This manner, however, would make the model highly sensitive to the data characteristics in specific scenarios, resulting in a lack of generalizability.
@@
Prediction-based. These methods aim to predict the ”normal” value of the current metric according to the adjacent observations. If the ground truth is far from the predicted value, an anomaly alarm will be triggered. The representative works include AR [44], LSTMAD [^2] As there is no source code available, we implement LSTMAD in two forms: LSTMAD-$\alpha$ in a seq2seq manner and LSTMAD-$\beta$ in a multi-step prediction manner. [45],
Reconstruction-based. These methods aim to denoise the anomaly points by the encoding-decoding phase. If the reconstructed value is far from the ground truth, it can be assumed that an anomaly occurs. The representative works include AE [46], EncDecAD [47], SRCNN [29], Anomaly Transformer [3], TFAD [28], and TranAD [48].
VAE-based. These methods hold the assumption that the distribution mapping and sampling process can enhance the denoising effect of the model and learn a more robust representation of normal patterns. The representative works include Donut [4] and FCVAE [49].
General Time Series Model. The designed universal model structure in this class can be used for all time series tasks including time series forecasting, classification, and anomaly detection. They tend to introduce a more general inductive bias that aims to better represent the general features of time series. The representative works include TimesNet [50], OFA [51] and FITS [52].
IV Experimental settings
IV-A Experimental Platform
IV-B Datasets
IV-C Learning Schema
IV-D Evaluation Metrics
Following the event-based evaluation criterion outlined in Sec. III-C, we establish two concrete metrics, namely $F1_{best}$ and $AUPRC$. $F1_{best}$ denotes the highest F1 score calculated under reduced-length point-adjustment when iterating over all possible thresholds. $AUPRC$ calculates the area under the precision-recall curve generated according to reduced-length point-adjustment, which is widely applied in cases where class imbalance is present in the data. While $F1_{best}$ can measure the best detection performance that the model can achieve on the current test set, $AUPRC$ provides a more nuanced evaluation of the model’s performance across different levels of recall, which can be important in anomaly detection. A model with a higher $AUPRC$ tends to be more robust.
IV-E Implementation details of Algorithms
V Experimental Results
@@
V-A Overall Performance
Figure 7: Overall performance ranking using different learning schema under reduced-length $F1_{best}$. Methods are ranked in descending order according to the average of all scores. The row names denote the names of methods, while the column names denote the names of the datasets. The best score of each column is underlined. Due to the inherent assumptions behind the statistical methods being incompatible with the all-in-one and zero-shot schemas, these methods are not included in the ranking, and the related score cells are set to grey without values.
@@ V-A Overall Performance
The overall ranking list is presented in Fig. 7, including the performance of each method using naive, all-in-one, and zero-shot schema, respectively. Since statistical methods, especially those based on streaming, heavily rely on the premise of data being identically distributed, these methods are not well-suited for the all-in-one and zero-shot learning schemas. Therefore, we only evaluate them under the naive schema. First and foremost, the performance of various models under the naive and all-in-one learning schema can exhibit substantial disparities (RI 1). For example, on the NEK dataset, although Donut performs well on the naive schema, its performance significantly degrades under the all-in-one schema. In contrast, FITS achieve better performance under all-in-one schema on this dataset. We list the impacts of different learning schemas on the performance of different methods in Table I. It is thought-provoking that these differences do not show a clear bias, suggesting that these variations are the result of a combination of multifaceted factors, which is discussed in detail in Sec. V-B. Additionally, given the consideration of errors arising from random dataset partitioning, the models demonstrated a relatively consistent performance across the all-in-one and zero-shot modes, as there may exist interdependence among time series in the same dataset (RI 2).
TABLE I: Performance differences on all datasets using different learning schemas. Best scores are highlighted in bold, and second-best scores are highlighted in bold and underlined. Performance improvements exceeding 5% are marked with $\uparrow$, while performance declines exceeding 5% are marked with $\downarrow$.
Method AIOPS NAB TODS WSD Yahoo UCR NEK
@@
Figure 8: Illustrations that some methods fail to effectively detect point-wise anomalies in certain scenarios. The black line represents the original curve, while the red line represents the anomaly scores provided by the method. Anomalous segments are highlighted in pink.
Generally, the statistical methods exhibit great performance on the dataset with relatively stable shapelets and low noise level (UCR), but work poor on other datasets with more noise. For deep learning based models, contrary to our expectations, the overall performances of the models that contain more complicated structures are not satisfactory (RI 3). We present a case (Fig. 8) from the AIOPS dataset to vividly demonstrate this phenomenon. Even though these point-wise anomalies seem to be quite evident, a significant portion of the methods are unable to handle such events effectively. As numerous factors can lead to a decrease in performance for DNNs, we summarize the major factors that result in the underperformance of some deep learning methods in anomaly detection tasks:
Poor noise resistance. The models’ anti-noise ability plays an important role in AD tasks. Compared to other domains like NLP or CV, time series data is affected by more factors during the collection process, resulting in more and intractable noise. Although models using simple structures (AR, FITS) sacrifices some feature expression capability, it’s hard for them to overfit noise. This contributes to the good results. For LSTM-series methods, due to the iterative encoding process, the forget gate in LSTMs is more likely to minimize the retention of low signal-to-noise ratio data in the hidden state. In constrast, due to the lack of special design for anti-noise, simply employing complicated backbones like transformer may lead to the overfitting of the noise, and finally result in the poor performance as shown in Fig. 8 (RI 4).
Lack of training data. As shown in Fig. 9, some models, especially the ones with complicated structures, achieve significant improvement in performance when using more training data. Models with complicated structures usually require a larger volume of training data to avoid overfitting due to their higher degrees of freedom and flexibility (RI 5). Similarly, models that employ intricate loss functions often necessitate a greater amount of training data, including diverse examples, to prevent the emergence of falling into local minima and even trivial solutions.
Figure 9: Case study on Yahoo_A3Benchmark-TS13 shows that lack of training data induces poor performance.
Trends with large periods. Some data may exhibit long-term trends or patterns that span large periods. These trend components can cause the data distribution to deviate from the comfort zone of the model, causing the model to calculate anomaly scores based on a biased data distribution it has never encountered before (RI 6).
Figure 10: Case study on Yahoo_A3Benchmark-TS14 shows that trends with large periods induce poor performance. The scores are obtained under the all-in-one learning schema, eliminating the impact of insufficient data.
Unreasonable inductive bias. Deep learning models rely on their architectural design and optimization algorithms to detect anomalies in a semi-supervised manner. If the chosen model architecture or optimization approach is not well-suited for point-wise anomaly detection, the models may have an unreasonable inductive bias that hampers their ability to accurately detect anomalies (RI 7).
Based on the above observations, it is imperative for newly proposed methods to place greater emphasis on these factors to avoid poor results on these distinct anomalies. We believe that the potential of the powerful representation capability of complex structures has not been fully explored yet in time series anomaly detection tasks. We strongly advocate prioritizing the refinement of module design or loss function in order to enhance the models’ resilience to noise and their capacity for generalization. Thus complicated backbones can better leverage their expressive power for temporal features without overfitting to noise.
V-B Performance in Fine-Grained Scenarios
@@
Merely referring to aggregated data at the granularity of datasets is insufficient to gain a profound understanding of the model’s performance. We carry out a fine-grained comparison of model performance from the perspectives of both learning schema and anomaly type. The UCR dataset is widely acknowledged for its comprehensive annotation of various anomaly types, making it a well-labeled dataset. Additionally, each time series in the UCR dataset is guaranteed to contain only one anomaly, which facilitates the categorization of different anomaly types. To further refine our analysis, we partition the UCR dataset into two subsets based on its supplemental materials [53]: one subset exclusively consisting of point-wise anomalies and another subset focusing solely on pattern-wise anomalies. We form a 2x2 matrix to reflect the variations in model performance under different conditions by incorporating the dimensions of ”naive” and ”all-in-one” learning schemas, along with different types of anomalies, as shown in Fig. 11.
@@
Figure 11: The 2x2 matrix reflecting the variations in model performance under different conditions. We denote prediction, reconstruction, VAE-based, and general time series methods in light red, yellow, green, and blue backgrounds.
@@
From the boxplots, no solution can always outperform the others in all situations (RI 8). It is also apparent that the majority of models exhibit a significantly higher detection performance for point-wise anomalies compared to pattern-wise anomalies regardless of the learning schema employed (RI 9). On a more detailed level, prediction-based models excel in identifying point-wise anomalies, as the steep peak/valley is unpredictable in most cases (RI 10). Due to the inclusion of a more diverse data distribution under all-in-one learning schema, the models are more likely to learn robust representations of patterns. As a result, under the all-in-one learning schema, there is a significant improvement in the detection performance of point-wise anomalies (RI 11). This observation is particularly pronounced in the case of the Yahoo dataset each time series of which is relatively short. With regard to pattern-wise anomalies which are considered to be more challenging, the situation becomes more complex and interesting. Methods that aim to generate the temporal window after projecting them to low-dimensional representations (Donut and FCVAE) surpass all others in detecting pattern-wise anomalies, while the performance greatly declines when under the all-in-one mode (RI 12). The former indicates that the assumption of low-dimensional representations hardly reconstructing high-level pattern-wise anomaly indeed works. An illustration is presented to confirm this in Fig. 12(a), where Donut can easily handle the ”smooth” high-level anomaly. However, the performance deteriorates or even becomes ineffective under the all-in-one mode (Fig. 12(b)). We hypothesize that the capacity of low-rank representations may struggle to cover the diverse data distributions present in various curves. Also, although the general time series methods (TimesNet, OFA, FITS) perform well in other time series tasks, they struggle to outperform methods specifically designed for anomaly detection tasks like LSTMAD and FCVAE (RI 13). This may indicate that there is a gap between time series anomaly detection tasks and general time representation tasks. For example, anomaly detection tasks may require models to have a stronger denoising effect. These gaps need special attention when designing models for anomaly detection tasks.
@@
Figure 12: Case study on CIMIS44AirTemperature2 which contains an easily understandable pattern-wise anomaly.
V-C Performance under Delay-Constraint
Issue d in Sec. III-C demands exceptionally high data set quality, whereas some datasets (including NAB, UCR, Yahoo) fail to offer labels without positional bias. Therefore, the result merely provide a approximate reference on these datasets. We manually set a relatively suitable latency limit ($K$) for each dataset based on the sampling frequency and label quality of the dataset to evaluate the performance of each method under strict real-time requirements. The overall results are shown in Table II. Compared to the results without delay-constraint, it can be seen that the performance of some methods significantly deteriorates in this setup, even approaching a 50% decrease (e.g., Anomaly Transformer on TODS). Although EncDecAD is inferior to FITS under all-in-one schema in Fig. 7, when considering delay-constraint, EncDecAD outperforms FITS on TODS, suggesting that some methods may have greater potential in terms of early detection capabilities.
TABLE II: Performance comparison using all-in-one learning schema under K-delay reduced-length $F1_{best}$ and $AUPRC$